# GOBench: Benchmarking Geometric Optics Generation and Understanding of MLLMs

### v1

@@ Abstract.
The rapid evolution of Multi-modality Large Language Models (MLLMs) is driving significant advancements in visual understanding and generation. Nevertheless, a comprehensive assessment of their capabilities, concerning the fine-grained physical principles especially in geometric optics, remains underexplored. To address this gap, we introduce GOBench, the first benchmark to systematically evaluate MLLMs’ ability across two tasks: 1) Generating Optically Authentic Imagery and 2) Understanding Underlying Optical Phenomena. We curates high-quality prompts of geometric optical scenarios and use MLLMs to construct GOBench-Gen-1k dataset. We then organize subjective experiments to assess the generated imagery based on Optical Authenticity, Aesthetic Quality, and Instruction Fidelity, revealing MLLMs’ generation flaws that violate optical principles. For the understanding task, we apply crafted evaluation instructions to test optical understanding ability of eleven prominent MLLMs. The experimental results demonstrate that current models face significant challenges in both optical generation and understanding. The top-performing generative model, GPT-4o-Image, cannot perfectly complete all generation tasks, and the best-performing MLLM model, Gemini-2.5Pro, attains a mere 37.35% accuracy in optical understanding. Database and codes are publicly available at https://github.com/Amber0614/GOBench.
@@ 1. Introduction
Geometric optics, encompassing optical phenomena governed by the Law of Rectilinear Propagation, the Law of Reflection, and Snell’s Law, is fundamental to visual appearance and composes most optical scenarios. (Loshin, 2015; Keating, 1988). However, achieving physically plausible optical generation (Tewari et al., 2020) and optical understanding (Zhao et al., 2024; Barron and Malik, 2012) is a significant challenge for MLLMs (Fu et al., 2024; Uddin et al., 2025; Li et al., 2024). Current models often produce visually appealing yet physically inaccurate results (Gani et al., 2023; Cherian et al., 2024; Abshari et al., 2024), with inconsistent lighting or incorrect optical effects (Zhang et al., 2025a), limiting their use in high-fidelity applications like realistic content creation or simulation (Hsu et al., 2024; Zhao et al., 2025; Zhang et al., 2025b; Zheng et al., 2025; Gu et al., 2024). This deficiency stems partly from a lack of benchmarks systematically probing these specific optical effects, unlike benchmarks for semantic understanding (Hu et al., 2021; Ma et al., 2024) or general VQA (Antol et al., 2015).
@@ 3.3. Results on Optical Understanding
The GOBench also evaluates the optical understanding and evaluative capabilities of 11 MLLMs, by comparing their assessment results of the GOBench-Gen-1k against human ground truth. Table 3.3 presents the performance using our primary metric, the Scaled Proximity Score (SPS) (Equation 1), with models sorted by their SPS in the Optical Authenticity dimension. The results indicate that even advanced MLLMs find optical understanding challenging. Gemini-2.5Pro (0506) leads in Optical Authenticity SPS with 37.35%, followed by Claude3-Opus (33.93%) and GPT-4o (latest) (32.63%). These top scores, considerably below 50%, highlight the substantial difficulty MLLMs in correctly judge the optical phenomena in the given images. Tops scores for aesthetic quality(32.03%) and instruction consistency(28.29%) further reveal the limitation of MLLMs’ ability to understand the geometric optical phenomena.

### v2

@@ Abstract.
The rapid evolution of Multi-modality Large Language Models (MLLMs) is driving significant advancements in visual understanding and generation. Nevertheless, a comprehensive assessment of their capabilities, concerning the fine-grained physical principles especially in geometric optics, remains underexplored. To address this gap, we introduce GOBench, the first benchmark to systematically evaluate MLLMs’ ability across two tasks: 1) Generating Optically Authentic Imagery and 2) Understanding Underlying Optical Phenomena. We curates high-quality prompts of geometric optical scenarios and use MLLMs to construct GOBench-Gen-1k dataset. We then organize subjective experiments to assess the generated imagery based on Optical Authenticity, Aesthetic Quality, and Instruction Fidelity, revealing MLLMs’ generation flaws that violate optical principles. For the understanding task, we apply crafted evaluation instructions to test optical understanding ability of eleven prominent MLLMs. The experimental results demonstrate that current models face significant challenges in both optical generation and understanding. The top-performing generative model, GPT-4o-Image, cannot perfectly complete all generation tasks, and the best-performing MLLM model, Gemini-2.5Pro, attains a mere 37.35% accuracy in optical understanding. Database and codes are publicly available at https://github.com/aiben-ch/GOBench.
@@ 1. Introduction
Geometric optics, encompassing optical phenomena governed by the Law of Rectilinear Propagation, the Law of Reflection, and Snell’s Law, is fundamental to visual appearance and composes most optical scenarios. (Loshin, 2015; Keating, 1988). However, achieving physically plausible optical generation (Tewari et al., 2020) and optical understanding (Zhao et al., 2024; Barron and Malik, 2012) is a significant challenge for MLLMs (Fu et al., 2024; Uddin et al., 2025; Li et al., 2024). Current models often produce visually appealing yet physically inaccurate results (Gani et al., 2023; Cherian et al., 2024; Abshari et al., 2024), with inconsistent lighting or incorrect optical effects (Zhang et al., 2025a), limiting their use in high-fidelity applications like realistic content creation or simulation (Hsu et al., 2024; Zhao et al., 2025; Zhang et al., 2025c; Zheng et al., 2025; Gu et al., 2024). This deficiency stems partly from a lack of benchmarks systematically probing these specific optical effects, unlike benchmarks for semantic understanding (Hu et al., 2021; Ma et al., 2024; Zhang et al., 2025b) or general VQA (Antol et al., 2015).
@@ 3.3. Results on Optical Understanding
The GOBench also evaluates the optical understanding and evaluative capabilities of 11 MLLMs, by comparing their assessment results of the GOBench-Gen-1k against human ground truth. Table 3 presents the performance using our primary metric, the Scaled Proximity Score (SPS) (Equation 1), with models sorted by their SPS in the Optical Authenticity dimension. The results indicate that even advanced MLLMs find optical understanding challenging. Gemini-2.5Pro (0506) leads in Optical Authenticity SPS with 37.35%, followed by Claude3-Opus (33.93%) and GPT-4o (latest) (32.63%). These top scores, considerably below 50%, highlight the substantial difficulty MLLMs in correctly judge the optical phenomena in the given images. Tops scores for aesthetic quality(32.03%) and instruction consistency(28.29%) further reveal the limitation of MLLMs’ ability to understand the geometric optical phenomena.
@@ 4. Conclusion
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grants No. 62225112 and U24A20220.